### v1

# Aligning LLMs for FL-free Program Repair

Abstract
Large language models (LLMs) have achieved decent results on automated program repair (APR). However, the next token prediction training objective of decoder-only LLMs (e.g., GPT-4) is misaligned with the masked span prediction objective of current infilling-style methods, which impedes LLMs from fully leveraging pre-trained knowledge for program repair. In addition, while some LLMs are capable of locating and repairing bugs end-to-end when using the related artifacts (e.g., test cases) as input, existing methods regard them as separate tasks and ask LLMs to generate patches at fixed locations. This restriction hinders LLMs from exploring potential patches beyond the given locations.
In this paper, we investigate a new approach to adapt LLMs to program repair. Our core insight is that LLM’s APR capability can be greatly improved by simply aligning the output to their training objective and allowing them to refine the whole program without first performing fault localization. Based on this insight, we designed D4C, a straightforward prompting framework for APR. D4C can repair 180 bugs correctly in Defects4J, with each patch being sampled only 10 times. This surpasses the SOTA APR methods with perfect fault localization by 10% and reduces the patch sampling number by 90%. Our findings reveal that (1) objective alignment is crucial for fully exploiting LLM’s pre-trained capability, and (2) replacing the traditional localize-then-repair workflow with direct debugging is more effective for LLM-based APR methods. Thus, we believe this paper introduces a new mindset for harnessing LLMs in APR.
I Introduction
Program repair is a critical part of the software cycle. To fix bugs in software systems, developers often need to dedicate substantial time (exceeding 35% of regular development time) for manual program repair [1]. To reduce such human effort, researchers have started exploring Automatic Program Repair (APR) methods [2, 3]. Based on how the patches are generated, APR methods can be categorized into heuristic-based [4, 5], constraint-based [6, 7], template-based [8, 9], and learning-based methods [10, 11]. Traditional methods often rely on pre-defined patterns for solving specific bug types (e.g., NullPointerException). Learning-based methods typically use large-scale, high-quality bug-fix pair data to train neural machine translation (NMT) models [12], which transform a buggy program to a fixed version [3]. Recently, as Large Language Models (LLM) [13] have shown strong code understanding abilities, many researchers have begun exploring LLM-based APR methods [14, 15]. With broad pre-training across diverse general corpus, LLMs can achieve impressive repair performance through prompt engineering [16] or minimal data fine-tuning [17].
Figure 1: First row: LLM sturctures and their training objectives. Second row: The training and inference objective is misaligned when using decoder-only LLMs for infilling-style APR. Third row: An intuitive way to align the gap: using LLMs for entire program completion rather than masked span prediction.
However, current application of LLMs is not well-adapted for program repair due to two major reasons. First, the objective of current approaches is misaligned with LLM’s training objective. Existing LLM-based methods [14, 18, 19, 15] typically adopt the infilling-style approach, i.e., an approach that predicts the masked span of sequence tokens given the preceding (prefix) and subsequent (suffix) code. These methods adopt the infilling style because its objective aligns with the objective of the encoder-only or encoder-decoder models used in their implementation. For example, AlphaRepair [14], one of the earliest LLM-based infilling methods, uses CodeBERT [20], an encoder-only model that is trained to denoise masked tokens. Recently, decoder-only LLMs (e.g., GPT-4 [21]) have shown stronger code comprehension capability [22], and recent research attempted to directly employ them as the patch generator for infilling-style APR [23, 24, 25], which demonstrate SOTA performance. However, as these decoder-only LLMs are trained to predict the next tokens given the past token context [13], there exists a misalignment between their training objective and the objective of infilling-style APR. Specifically, while decoder-only LLMs are proficient at generating a complete program because of the sufficient code corpus in pre-training, we hypothesize that these LLMs are less proficient in patch generation on masked buggy code due to the sparsity of parallel <masked, denoised> code corpus in pre-training. As noted in recent studies about LLM, objective misalignment can yield significant sub-optimal performance [26, 27].
Second, the current workflow limits LLM from fully exploiting its pre-trained capability. Existing approaches still adhere to the workflow of first using fault localization (FL) tools [28] to obtain a ranked list of potential fault locations, and then using APR methods to generate patches by modifying the buggy program at these locations [2, 3]. However, LLMs already have the capability to fix trivial syntax bugs in their generated code independently by referring to different artifacts (e.g., error report from compilers and virtual machines) [15, 29, 30], illustrating the capability of identifying bugs without FL tools. Considering LLM’s superior code comprehension ability, and several studies show that current FL tools may fail to provide precise fault locations [31, 32], asking LLMs to generate patches at provided locations might contrarily hinder them from exploring broader potential patch space beyond the provided ones. A recent study also revealed that LLMs fail to make good use of the given buggy lines and tend to over-rely on them [17], whereas another study suggested using a more flexible fault localization to enhance APR [33]. As observed in these studies, the traditional localize-then-repair workflow may lead to ineffective LLM-based APR.
In this work, we explore a novel way to use LLM for program repair. We propose that (1) aligning the output from infilling discrete hunks to completing entire functions can better attain the training objective, and (2) allowing LLM to locate and repair bugs with artifacts in a human-like manner can further improve its APR performance. Based on these insights, we implemented D4C (i.e., Direct Debug Drives Decent Code), a straightforward end-to-end APR approach without complex prompting or extra training. We conducted experiments on a total of 1027 bugs associated with individual functions (i.e., single-function bugs) from Defects4J [34] and DebugBench [35]. D4C repaired 180 out of the 437 single-function bugs in Defects4J, outperforming the state-of-the-art (SOTA) APR methods with perfect FL [24] by 10%. Meanwhile, each patch in D4C was sampled only 10 times, accounting for 10% used by SOTA methods (which use 100–5000 samples). Our results show that by fully aligning the output format and augmenting the input prompt, D4C can achieve the best end-to-end APR performance without FL, additional fine-tuning, or multi-round dialogue. This paper is not aimed at proposing D4C as a new APR technique, but rather to introduce a new mindset or paradigm for better harnessing LLMs in APR in the future.
In summary, this paper makes the following contributions.
Problem reformulation:
We reformulate infiling-style repair problem as a program refinement problem. Our experiments show that asking LLMs to generate an entire refined function can align with the pre-training objective of decoder-only LLMs, leading to SOTA APR performance.
New APR workflow:
Instead of relying on fault localization to obtain the list of locations to modify as in traditional APR workflow, we show that allowing LLMs to end-to-end locate and repair bugs by using diverse types of artifacts (e.g., failed tests, error messages, and code comments) in a human-like manner can further improve its APR performance.
Implementation and evaluation:
We implemented D4C based on these insights, and the evaluation results on two widely used benchmarks (i.e., Defects4J and DebugBench) have demonstrated its effectiveness.
II Background and Motivation
II-A LLM Architectures and Training Objectives
Background: A large language model (LLM) is a language model consisting of a neural network with many parameters (typically billions of weights or more) trained on large quantities of unlabelled corpus using self-supervised learning [36]. The LLMs usually adopt the Transformer [37] architecture or its sub-structures (i.e., encoder or decoder). Based on their architectures, LLMs can be categorized into three types: encoder-only, decoder-only, and encoder-decoder LLMs. LLMs in the same architecture also share the same or similar training objectives. The details of the architecture and corresponding training objectives are shown in the Fig. 1.
Encoder-only LLMs, such as BERT [38] and its variants like CodeBERT [20], have a bidirectional transformer encoder structure. They are typically trained on the masked language modeling objective (i.e., MLM), aiming to denoise and reconstruct the masked tokens via understanding the surrounding context. As shown in Eq. 1, the loss of MLM training objective can be explicitly represented as the log-sum of the conditional probabilities of generating the masked token when all the unmasked tokens are known.
$\mathcal{L}_{MLM}=-\frac{1}{M}\sum_{i\in M}\log P(\hat{t}_{i}=t_{i}|t_{k\notin mask};\theta)$ (1)
where $M$ is the total number of masked tokens.
Decoder-only LLMs, including GPT series [13, 21] and LLaMA series [39], have an autoregressive transformer decoder structure. They are mainly trained on the causal language modeling objective (i.e., CLM), aiming to predict and complete the next tokens via following the prefix context. As shown in Eq. 2, the loss of CLM training objective can be explicitly represented as the log-sum of the conditional probabilities of generating the next token when all the preceeding tokens are known.
$\mathcal{L}_{CLM}=-\frac{1}{N}\sum_{i\in N}\log P(\hat{t}_{i}=t_{i}|t_{1},t_{2},...,t_{i-1};\theta)$ (2)
where N is the total number of the input tokens.
Encoder-decoder LLMs, such as T5 [40] and its variants like CodeT5 [41], have a complete transformer structure. They use an encoder to embed the input text into vector representations, and a decoder to causally generate new tokens after the input prompt. Specifically, T5 series LLMs are often pre-trained on the denoising objective, which aims to generate reconstructed spans (i.e., a sequence of adjacent tokens) for the masked span in the prompt. As shown in Eq. 3, the loss of denoising training objective can be explicitly represented as the log-sum of the conditional probabilities of generating the masked token when all input tokens are known.
$\mathcal{L}_{Denoising}=-\frac{1}{M}\sum_{i\in M}\log P(\hat{t}_{i}=t_{i}|t_{k\in input};\theta)$ (3)
where M is the total number of the masked tokens. While this process is similar to the MLM objective of encoder-only LLMs, it only generates the reconstructed tokens from the masked position, ignoring other parts (e.g., “I” and “flowers” in Encoder-Decoder LLM in Fig. 1). Similar to decoder-only LLMs, it also has an unfixed length of text generation.
Motivation: As introduced, both the encoder-only and encoder-decoder (especially T5 series) LLMs involve the corruption and denoising process during pre-training, which substitutes some tokens in the input text for mask tokens, and reconstruct these tokens from the mask tokens. To make effective use of them for APR, Xia et al. [14]. proposed the infilling-style APR. It aims to replace the entire input buggy hunk with a mask token (e.g., <INFILL>) and prompts the model for direct inference, restoring the fixed hunk from the mask. The infilling-style APR can be implemented differently across various models. For instance, in CodeBERT, it recovers the correct code from masked hunks while leaving the rest of the input tokens the same [20]. Meanwhile, in CodeT5, it only outputs the correct code restored from the masked hunks, excluding all other input [41]. These task modeling strategies align with their model training objectives.
Figure 2: An example of current automated program repair workflow.
Figure 3: An example of two different APR paradigms. First row: Locate and repair subsequently may cause many invalid attempts when the localization is wrong. Second row: Locate and repair simultaneously can mitigate the limitation of patching at specific locations. (The wavy line denotes the fault location)
Recently, several APR approaches use decoder-only LLMs as they have exhibited a strong capacity for program comprehension and generation. Influenced by the prior use of infilling-style approaches, it has been generally assumed that such methods would be useful for APR. Thus, prior techniques use decoder-only LLMs to generate discrete fixed hunks for the given buggy hunks, as illustrated in Fig. 2. However, except for very few models (e.g., InCoder [42]), decoder-only LLMs barely incorporate denoising objective during pre-training. As the <masked, denoised> parallel corpus is rarely available, it becomes challenging for the models to learn the generation of the fixed hunk from the masked tokens without further adaptation after the next token prediction pre-training. Such an objective misalignment might hamper LLM’s performance on specific tasks extensively, such as the text classification task [27] and the QA task [26], whose task objectives are misaligned with the training objective of decoder-only LLMs. Thus, we have the following insight.
Insight 1: Modifying decoder-only LLMs’ output from discrete fixed hunks to the entire refined program can better align the inference objective to the training objective, thus significantly enhancing APR performance.
This insight was inspired by previous approaches focusing on optimizing model performance on specific tasks by aligning the pre-training objectives of the model [27, 26]. We will verify our hypothesis in Sec. IV-C via comparing completion perplexity of two output formats (i.e., fixed hunks or complete function) on white box LLMs (i.e., Mixtral-MoE).
II-B APR Techniques and Workflow
Background: Given a buggy program and an artifact (e.g., usually a test suite with at least one failed test), automated program repair (APR) approaches generate a fixed program that fulfills a correctness criteria (e.g., passing all tests). As illustrated in Fig. 2, a traditional APR workflow typically encompasses three steps: (1) fault localization, (2) patch generation, and (3) patch validation. Particularly, APR research mainly focuses on the patch generation step and obtains the fault locations using existing FL techniques (e.g., statistical fault localization) or uses perfect FL results in evaluation.
Based on their patch generation strategies, APR methods can be categorized into heuristic-based [4, 5, 43, 44, 45, 46, 47], constraint-based [7, 48, 49], template-based [50, 8, 9], and learning-based  [51, 10, 11, 52, 53, 54, 55]. The non-learning-based approaches are usually restricted by a limited set of program transformations, causing the generation of a large number of invalid patches. To identify location to apply these transformations, the current strategy is to first generate a ranked list of suspicious fault locations using FL tools. The APR tool then sequentially generates patches for each provided fault location. Each of the generated patches is then validated using the given artifacts (e.g., test cases). With the advent of deep learning (DL), learning-based APR tools based on neural machine translation (NMT) have emerged. These tools outperform the traditional methods [3] via learning code semantics on a large bug-fix parallel training corpus without specially designed patch templates nor search heuristics. Moreover, since NMT models generally have a limited parameter size, their training costs and inference efficiency are manageable. Therefore, they are well-suited for the traditional APR workflow.
Figure 4: The workflow of D4C. It uses the buggy code and its corresponding documents, failed tests, and test info (e.g., error message) to construct the prompt for one-shot prompting-based program repair without a specific fault location (usually provided by FL tools).
Motivation: In recent years, there is a growing interest in using LLMs (e.g., GPT-4 [21]) for APR [14, 56, 33, 54]. Researchers usually follow the current APR workflow, positioning LLM as a new patch generator to replace former NMT models. However, taking LLM as a simple substitute in the patch generation step in the current workflow is an under-utilization of its pre-trained knowledge, since these models exhibit near-human capabilities in complex problem analysis and reasoning. For example, SELF-DEBUGGING has been proposed with the concept of allowing LLMs to progressively refine their generated code by engaging in multiple self-dialogue rounds using question explanations and execution traces, eventually producing high-quality code [30]. As shown in Fig. 3, since FL tools may not always provide perfect predictions of fault locations, restricting LLM to fixing the provided locations may lead to time waste in validating patches at incorrect locations. Moreover, as LLM’s inference overhead and time cost is much higher than traditional APR tools, using LLM to generate patches for all given locations may result in huge resource waste. Thus, we have the insight below:
Insight 2: Prompting LLMs with buggy programs and corresponding artifacts (e.g., failing tests) can enable them to locate and repair bugs simultaneously without FL, thus further improving their APR performance.
This insight arises from the APR methodology of programmers who typically identify bugs through tests and documents, and subsequently fix bugs based on these test results. Moreover, existing research revealed that models can produce decent fixes on basic compile bugs when supplemented with compiler errors [15], and their patch generation capability can be further improved with the assistant of the given failed test information [24, 30]. Thus, we believe this approach can overcome the limitation of generating patches within a limited number of locations, thereby further enhancing LLM’s APR performance. We will verify our hypothesis in Sec. IV-C via comparing the number of correct patches of two input formats (i.e., w/ or w/o artifacts) on Mixtral-MoE and GPT-4.
III D4C: Direct Debug Drives Decent Code
Based on these two insights, we developed an APR framework D4C, i.e., direct debug drives decent code. As shown in Fig. 4, when presented with a new buggy program [^1] In our implementation, the buggy program refer to a single function, due to the limitation of LLMs’ context window length., D4C uses the related artifacts, including the documents and failed test information, to construct a bug report. Then, D4C will use this bug report to instruct the model to generate a refined version of the buggy program. To make LLM follow D4C’s input-output format, we adopt a one-shot prompting strategy by providing a handcrafted input example and its output label. D4C does not require any additional fine-tuning. It only changes the output format to align with the pre-training objectives, and uses test information and documents (e.g., Javadoc comments) in the input prompt to enable repair without the prior knowledge of the correct fault locations. We implemented our approach on GPT-4 series models and Mixtral-MoE models. Next, we will introduce our problem definition (Sec. III-A) and the design details of D4C, including model selection (Sec. III-B), artifact processing (Sec. III-C), prompt construction (Sec. III-D), patch generation (Sec. III-E) and validation (Sec. III-F).
III-A Problem Definition
We model the APR task as a completion task that aims at generating a complete refined program based on a buggy program and associated artifacts. This is different from previous LLM-based APR approaches, which treated the APR task as an infilling or cloze task [14]. Specifically, the optimization objective of D4C can be formally written as:
${\arg\max}_{\theta}\ \ P(t_{\text{Fixed Program}}|t_{\text{Buggy Program}},t_{\text{Artifacts}};\theta)$ (4)
and the objective of existing infilling-style APR is:
${\arg\max}_{\theta}\ \ P(t_{\text{Fixed Hunk}}|t_{\text{Buggy Program}};\theta)$ (5)
where $t$ means the tokens, $\theta$ is the trainable parameters of LLMs. The reason for the difference between D4C and the previous methods in task modeling is that D4C is designed to exploit the pre-training capability of decoder-only LLMs trained on CLM objective, whereas infilling-style APR is designed to exploit the pre-training capability of encoder-only LLMs trained on MLM objective or encoder-decoder LLM trained on denoising objective.
III-B Model Selection
The LLM itself is always the most vital part of LLM-based applications, which determines the upper bound of the capability for handling specific downstream tasks. Since the goal of proposing D4C is to illustrate an adaptive way to use current advanced decoder-only LLMs for APR, the backbone of D4C should be a decoder-only LLM which has been pre-trained on a substantial code corpus via next token prediction objective. Moreover, D4C is not a complex approach aiming at teaching LLMs to perform APR. Instead, its APR ability primarily comes from aligning the model’s completion format with its pre-training corpus format, e.g., a complete function. Thus, to enable D4C exhibit the state-of-the-art APR performance, we choose the most advanced LLMs (e.g., GPT-4 and Mixtral-MoE) to serve as the backbone in our implementation, as they have long context windows for generating refined version of most of the buggy functions. We will not dive into their architecture or pre-training details in this paper, but it is worth noting that the backbone for D4C can be changed to other LLMs with next token prediction (CLM) training objective. To best ensure our replicability, we employed fixed remote API checkpoints (e.g., gpt-4-0613) or fixed local model versions (e.g., mixtral-moe-instruct-v0.1). For reference, we also provide comparative experiments about using different backbones in Sec. IV-C.
III-C Artifact Processing
Before patch generation, D4C needs to extract related artifacts for subsequent prompting. As shown in Fig. 4, D4C draws (1) program documents or comments that describe the general purpose of the code and its input-output data types, (2) the inputs and expected outputs of the failed test cases, and (3) error messages from those failed test cases, corresponding to the buggy code. Our underlying motivation for providing these artifacts to the LLM is that LLM has gradually exhibited human-like analysis and reasoning abilities recently [29]. Instead of treating LLM as a tool as in the traditional mindset of training NMT models from scratch, we should treat it as an intelligent agent. From this perspective, we contemplate how humans debug, and present this procedure to LLM for execution. Specifically, in the human program repair process, we need to first intuitively understand the function’s purpose so we refer to related code documents or function-level comments. Then, we wonder what input can reproduce the bug. Thus, we will check the relevant failed test cases and error messages. Finally, in the absence of communication or confirmation from other developers (i.e., in a rubber duck debugging [57] cases), we can only use these artifacts to simultaneously locate and repair the bug. Therefore, we extract artifacts and then construct bug report for prompting LLMs to repair buggy program.
In specific deployment scenarios, we extract each buggy program’s test results and documentation through regex matching, and store them in different fields of a table, categorized by the bug ID. It is noteworthy that the implementations of the artifacts processing procedure may slightly differ due to various data characteristics. For instance, although we default to using function-level comments as documents, if certain functions do not have this artifact, D4C can alternatively exploit related README documents. Moreover, if some artifacts are unavailable in some scenarios (e.g., the test cases of the automated testing tools are unknown), D4C will replace these in the table with a placeholder statement, like “This program does not possess any known test cases.
III-D Prompt Construction
Figure 5: The prompt structure of D4C. The details of the code are omitted.
After the extraction of artifacts, D4C assembles these artifacts into a comprehensive bug report. As shown in Fig. 4, the report template is concise, with only the most basic notations of the different components in the bug report (e.g., program and comment), without providing any complex instructions or requirements. This allows LLM to focus more on the bug report, thus accomplishing our goal of inducing and exploiting LLM’s pre-trained knowledge and capabilities for program repair, rather than “teaching” the model to repair. Following this, D4C integrates the bug report into a complete prompt, which is carefully designed to include a concise role-play system instruction (e.g., You are an AI debugger …) and a pre-determined example consisting of a handcrafted bug report and its refined program. Our underlying goal for providing the pre-determined example to the LLM is to stimulate LLM’s in-context learning ability [13], i.e., synthesizing the parallel report-refinement pair in the prompt to further combine the instruction to understand the objective of the APR task and the mapping relation between input and output. This one-shot in-context learning approach also allows us to restrict the output format of the LLM to a refined function instead of generating other output formats related to APR tasks (e.g., unit diff patch [58]). Moreover, in-context learning assists us in automatically extracting the corrected functions from LLM response for automated patch verification. Specifically, as illustrated in Fig. 5, we enhance the output of the example with markdown-style code formatting control symbols (e.g., ‘‘‘java). This enables us to perform regex matching based on these control symbols when extracting the function output for subsequent processes. We do not add more than one example in the prompt, since our goal is to make LLMs follow our designed output format, thereby verifying our insights, rather than using complex prompting strategy to design a powerful APR method. The pre-determined instruction and the example remain fixed and unchanged in the inference stage.
[…]
III-E Patch Generation
Unlike traditional APR tools that generate patches specifically for locations that need modification, the patch generation step of D4C refines the entire program in the inference stage. This is primarily because the core goal of D4C is to explore the extent of APR enhancement by aligning the inference objective to decoder-only LLM’s pre-training objective. Furthermore, we do not introduce an additional patch ranking method, which is widely adopted by current APR approaches. This is because D4C only needs to generate at most 10 samples during decoding (i.e., beam search), which is significantly less than the sampling number of existing methods (i.e., over hundreds of sampling). Thus, D4C can subsequently verify all patches instead of the time-consuming process of verifying hundreds of patches. Additionally, although D4C can rank the patches by comparing their naturalness (quantified by perplexity or entropy) [59, 60], a prior study indicates that it is inaccurate to identify the correctness of the program using naturalness [61]. Therefore, we do not use an additional re-ranking strategy to save the computational cost incurred by the perplexity calculation. Similar to most LLM-based techniques that focus on method-level generation [62], we assume that the buggy function is provided where D4C can select lines within the function to modify. In a real-world scenario if the buggy function is not given, we can refine candidate buggy functions by iterating through all functions executed by the failed tests.
III-F Patch Validation
After generating the candidate patches, D4C extracts the refined program and uses them to replace the buggy program in the original source code. Then, D4C runs the corresponding test suite to find plausible patches that compile successfully and pass all tests. Following prior approaches [14, 63, 24], we then manually check the correctness of the plausible patches to find the correct patch (i.e., a patch that is exactly the same or semantically equivalent to the human-written patch).
IV Evaluation
This section aims to answer the research questions below:
  • RQ1: How does D4C compare against LLM-based APR methods? We compared D4C to state-of-the-art APR methods with perfect FL. We aim to investigate the improvement from aligning output and enhancing input as mentioned in Sec. I.
  • RQ2: How effective are the two insights that drive the design of D4C? We aim to show that modifying the output to a refined function can better align the training objective, and prompting the bug report to locate and repair bugs without FL can enhance APR performance.
  • RQ3: How do different components and parameter affects D4C? We conduct an ablation study and a sensitivity analysis on D4C. We aim to quantify the contribution of each prompt component and characterize the performance of D4C in different parameter settings.
IV-A Experiment Setup
IV-A1 Environment and Implementation
We use a black-box LLM (i.e., gpt-4-0613) and a white-box LLM (i.e., Mixtral-8x7B-Instruct-v0.1) via OpenAI APIs [64] and 8xA100 NVIDIA GPU server for program repair. We use Python 3.9 to implement the inference and evaluation scripts in a local machine with Ubuntu 20.04.5 LTS. To enhance LLM’s instruction following ability, we manually crafted a bug-fix example. This example is employed in each experiment to construct 1-shot prompting. To adapt to different experiments, we modified its input-output format according to the corresponding experiment requirement. Following the previous work, we use the default temperature=1.0. However, we adopt a much smaller sampling number of 10, as the inference budgets (fee, time, etc.) for LLMs are much larger than traditional methods. Unless otherwise specified, all experiments adhered to this setting.
IV-A2 Datasets
Our experiments are conducted on 1027 logic single-function bugs from DebugBench (590) and bugs from Defects4J (437). Specifically, DebugBench [35] is a new debug benchmark designed to counter data leakage (by implanting bugs into source data with GPT-4), which contains a total number of 4,253 bugs from Java, C++, and Python3 from LeetCode. As this paper focuses on APR (instead of syntax error repair), we only evaluate the logic bug type in DebugBench. Since DebugBench is not exposed to the threat of data leakage, we use it for our insight validation (RQ2). However, due to the complexity to reproduce current baselines on this new benchmark, we do not use it for the comparison between D4C and baselines. Defects4J [34] is a widely used APR benchmark with 835 real-world bugs from 17 open-source repositories. Following previous work [23, 24, 25], we separate the single-function bugs of Defects4J to v1.2 (203 bugs) and v2.0 (234 bugs). Since it contains data from previous versions of open-source projects where most code models have been trained on, it faces the threat of data leakage. However, this threat is believed to be less severe in APR compared to other domains [17] due to the sparsity of bug-fix parallel corpus in the training data, meaning that model can only learn the individual buggy or fixed version of the program, not their pairs. Despite this, the potential threat of data leakage could still undermine the validity of hypothesis verification. Hence, Defects4J is mainly used as a reference for comparing D4C with existing baselines. The statistic of the benchmarks are shown in Table I.
TABLE I: Statistics of the benchmarks: The number of single-funciton bugs and the threat of data leakage
Benchmark Bug number Data leakage Language
Defects4j 437 Yes Java
DebugBench 200/194/196 No C++/Java/Python3
IV-A3 Metrics
Following the previous work, we use the number of correct patch to evaluate the APR effectiveness. Specifically, if a patch can pass all unit tests, then it will considered as a plausible patch. If this plausible patch truly resolves the bug, rather than overfit to pass the unit test only (e.g., via referring the provided failed tests in the prompt), then it will be confirmed as a correct patch. In our evaluation, we adhere to prior studies [14, 24, 63] and manually identify whether a plausible patch truly resolves the bug. Notably, the bugs from DebugBench are collected from LeetCode, and the LeetCode programs that can pass all the unseen test suites from the website are deemed as correct programs. In our experiment, we can only access the fixed test examples of each bug in our bug report, rather than the failed unseen test from LeetCode. Thus, we also determine the patches that can pass all the LeetCode tests as correct. To distinguish from manually checked correct patches, we refer to patches can pass the LeetCode validation as verified patches. More details of correctness validation are discussed in Sec. V.
IV-A4 Baselines
In RQ1, we selected several state-of-the-art LLM-based APR methods as baselines, including AlphaRepair [14], FitRepair [19], ChatRepair [24], RAP-Gen [18], and Repilot [63]. We did not compare D4C with learning-based APR methods trained from scratch since our goal is to illustrate the effectiveness of D4C among all the LLM-based approaches. All these baselines employ the infilling objective. In RQ2, we chose different input and output format for comparison to show the superiority of aligning output format and providing artifacts in the input. We also chose prompts that add or delete different information (e.g., function comments) as baselines to investigate the contributions of different information to end-to-end APR.
TABLE II: The number of correct patches generated by different approaches on Defect4J.
Method Model v1.2 v2.0 Sum Sampled
AlphaRepair CodeBERT 52 34 86 5000
Repilot InCoder 66 50 116 5000
RAP-Gen CodeT5 72 53 125 100
FitRepair CodeT5 89 44 133 5000
ChatRepair GPT-4 114 48 162 100-200
D4C GPT-4 84 96 180 10
IV-B RQ1: Comparison against LLM-based APR
We compare D4C with five state-of-the-art LLM-based APR approaches on Defects4J, the most commonly used benchmark for APR. Since some of the baselines are not open-sourced, we cannot re-run them to reproduce the result. Thus, we reuse their Defects4J results reported in the original paper. Although using Defects4J may have the the risk of exposing its data to LLM’s training corpus, we still use it for evaluation because (1) existing work [17] has shown that data leakage is less of a concern for APR compared to other code-related tasks as the training corpus often contains at most the individual versions of the program (i.e., learning buggy or fixed version only), and LLMs can hardly learn the APR tasks without bug-fix pair corpus during training; and (2) the baselines have also been evaluated on Defects4J (hence, they are subject to the same data leakage risk). Following prior work, we retain this setting to ensure fairness of the comparison.
The comparison result is shown in Table II. We calculated the number of correct patches on Defects4J (v1.2, v2.0, and their sum), and the number of patches sampled per bug (notated as “Sampled”). D4C largely outperforms all the existing state-of-the-art methods provided with perfect bug locations for infilling-style APR. Specifically, D4C (GPT-4) can generate almost 10% more correct patches than ChatRepair, the latest APR methods which uses GPT-4 for multi-run dialogue with provided perfect bug locations. Additionally, D4C does not require FL tools to provide perfect bug locations, which means that D4C is a more practical tool in development, since perfect FL is difficult, if not impossible, to achieve in practice. Moreover, D4C only needs to sample 10 times for each bug in patch generation, which is 90% fewer than that required by the most efficient baseline (sampling at least 100 patches). The fewer generated patches also illustrate the efficiency of D4C in fast program repair. Therefore, the results suggest D4C is an effective and efficient APR approach. However, due to the potential of data leakage, we cannot directly verify our insight of D4C using this comparison. Thus, we conduct more in-depth experiments on DebugBench in Sec. IV-C and Sec. IV-D.
IV-C RQ2: Effectiveness of the two insights
RQ2 aims to validate the effectiveness of the two insights introduced in Sec. II, and how the two insights contribute to the overall performance improvement in D4C. We design two additional questions for the insights: (IQ1) Does changing output to complete function align the training objectives and result in better performance? (IQ2) Does providing artifacts to the model enable an end-to-end APR and achieves a better performance? Specifically, we evaluate the two insights on DebugBench since there is no threat of data leakage. Thus, we use verified patches (i.e., patches verified by LeetCode unseen tests) in DebugBench for the evaluation.
Figure 6: Examples of input and output format used in RQ2.
IV-C1 Validation of Insight 1
To answer the IQ1 question, we calculate the average perplexity (i.e., measures how well a model predicts samples) for various output format on the white-box model (i.e., Mixtral-MoE). Lower perplexity values indicate the model is more confident at predicting a given sequence. Specifically, we measure perplexity by calculating the loss of CLM objective (Eq. 2). Measuring perplexity is the most straightforward approach to evaluate whether the generated content aligns with the distribution of training corpus, as the LLM is trained to minimize the CLM objective. Specifically, we adopt two output formats, including the fixed hunk patches that is used by infilling-style APR (denoted as Hunk) and the full-function output used by D4C (Func). Fig. 6 shows the examples of these output format. In our implementation, we also carefully modified the system instructions to adapt to each output setting. To control variables, we also calculate the perplexity of generated output (O), and the entire input&output (IO). Thus, we can compare the perplexity among different output format to validate our insight, ignoring the difference between each input format.
TABLE III: Comparison of different settings of D4C on DebugBench. We measure perplexity and verified patches to validate insights.
Format Perplexity Mixtral (#Verified) GPT-4 (#Verified)
O IO C++ Java Python C++ Java Python
Mask-Hunk 8.59 2.68 66 39 63 140 139 129
Mask-Func 3.01 1.58 99 91 76 160 153 147
Report-Hunk 8.50 2.64 72 64 67 144 140 133
Report-Func 1.39 1.79 118 125 104 177 163 161
Table III shows the results. We observe that when the output is an entire function, the output perplexity is much lower than using the fixed hunks as the output, regardless of the input. This directly validates that generating a complete function is the better way to align the training objective. To further validate that it is the objective alignment that helps D4C achieve a better APR performance, we evaluate the verified patches generated in each case. The result illustrates that using complete function as output can result in generation of more verified patches with lower perplexity than using discrete hunks. Moreover, when only generating the fixed hunks as the output, the output perplexity is much higher than its perplexity calculated with the whole input and output sequence. This indicates that completing discrete hunks only is misaligned with LLM’s training objective. Overall, these findings illustrate the effectiveness of our first insight.
IV-C2 Validation of Insight 2
To answer the IQ2 question, we evaluate the patch generation ability on DebugBench using two different input formats, i.e., the bug report (notated as Report) and the buggy program where the buggy hunks are masked (Mask). To control variables, we also conducted experiments using the two previously established output formats as shown in Fig. 6. For the generated patches by Report-Hunk, as we do not provide a specific location to modify and cannot precisely determine where the generated patch should be applied, we instruct Mixtral-MoE to help us to automatically apply the generated patches to the original buggy program before patch validation. Table III shows the number of generated verified patches on both Mixtral-MoE and GPT-4.
The results indicate that Report-Func is the most effective among all evaluated groups, including Mask-Hunk, which is the input-output format of infilling-style APR. This phenomenon aligns with prior evaluations on Defects4J (RQ1). It is worthwhile to note in the experiment in Sec. IV-B, the effectiveness of D4C stems not only from the bug report design but also the objective alignment of the output. However, even when we use the same output format in this experiment, the performance of using bug report as input still outperforms that of using program where the buggy locations are masked. This suggests that bug report helps not only in fault localization but also give more information for patch generation. Thus, this phenomenon validates our second insight.
IV-D RQ3: Ablation study and sensitive analysis
TABLE IV: The number of patches generated for our ablation study and sensitive analysis. We measure plausible patches for Defect4J and patches verified via unseen tests for DebugBench.
Benchmark Defects4J (#Plausible) DebugBench (#Verified)
v1.2 v2.0 Sum C++ Java Python Sum
w/o Document 93 99 192 175 162 158 495
w/o Test 80 89 169 173 159 159 491
w/o Message 81 93 174 169 162 161 492
- Mask 53 73 126 160 153 147 460
Samp=1 60 73 133 163 151 145 459
Samp=3 91 88 179 177 163 161 501
Samp=10 (default) 96 114 210 181 171 169 521
Temp=0.0 61 75 136 165 154 146 465
Temp=1.0 (default) 96 114 210 181 171 169 521
In this section, we aim to quantify the contribution of each component, and investigate the sensitivity of D4C to the sampling number and temperature. We do not additionally validate the sensitivity on different prompts since we have carefully investigated the impact of different prompt formats in RQ2. We use GPT-4 as the backbone, set the default temperature to 1.0, and use the sampling number of 10 by default. As the manual validation of patch correctness on Defects4J is time-consuming, we use the plausible patches in this section. We also use verified patches to evaluate the repair performance on DebugBench. We design the baselines below to check the effectiveness of each component:
w/o Document: D4C without including the program documents or function comments in the prompt.
w/o Test: D4C without including inputs and expected outputs of the failed test cases in the prompt.
w/o Message: D4C without including error message from the failed tests in the prompt.
- Mask: D4C using masks to identify bugs without including any artifacts in the prompt (Mask-Func).
Table IV shows the results of our ablation study and sensitive analysis. Overall, the results show that each component is important in guiding D4C in generating plausible/verified patches across the two benchmarks. We also observe that the “- Mask” baseline are the least effective because D4C cannot use any information from the artifact to fix the defect. Meanwhile, among the first three baselines where we remove one information from the prompt, we notice that program documents (w/o Document) are the least effective as D4C is able to generate the greatest number of plausible/verified patches without using the document. This is expected as the error message and test are more closely related to the defect than the comments that merely describe general code features.
In terms of sample size, we notice that a sample size of 10 is the optimal setting to allow D4C to find the correct patches. We found that increasing the sampling number from 1 to 3 significantly increased the number of plausible and verified patches, but further increasing the sampling number to 10 resulted in a smaller improvement. Since D4C can already achieve satisfactory results when setting the sampling number to 10, we use it as the default sample number for D4C. We also observed that setting the temperature to 0 results in fewer plausible and verified patches. Thus, we adopt a multi-sampling for D4C rather than greedy decoding (temp=0).
V Discussion
V-A Threats to Validity
External Threats. During the manual patch correctness analysis, we follow the procedure of previous work [14, 24, 19, 63]. To mitigate the potential bias in the manual analysis, two authors independently confirm the patch correctness. Any patches with disagreement were presented to the third author for review. However, since we could not submit the patch to the developers of the corresponding repository for confirmation, this threat could not be completely eliminated. We also release our results for public verification. As the effectiveness of the model may vary in different settings, our results may not generalize beyond the studied settings and other programming languages beyond the supported ones in Defect4J and DebugBench. We mitigate this threat by reusing configurations in prior work, and widely used benchmarks.
Internal Threats. One threat to internal validity lies in the correctness of the implementation and scripts that can affect our reported results. To mitigate this threat, we have carefully reviewed all the source code and scripts. Another internal threat comes from the potential data leakage problem (i.e., using Defects4j in RQ1, which may have been incorporated in LLM pre-training). We have discussed this threat in RQ1 (i.e., data leakage is not significant on APR [17], and the baselines have also been evaluated on Defects4J). To eliminate this threat in our insight validation, we selected DebugBench, a latest, leakage-free benchmark, for evaluation. This not only keeps a fair comparison in RQ1, but also ensures the credibility of our conclusions in RQ2 and RQ3.
Figure 7: A plausible patch generated at developer patch locations (Jsoup-19).
V-B Qualitative Analysis of Plausible Patches
Figure 8: A correct patch generated at non-developer patch locations (JxPath-8).
Table IV shows that D4C generates a total of 210 plausible patches but 30 are incorrect with respect to the human patches. Fig. 7 shows one example of the incorrect patch. Since the failed test contains the "src" tag as input, LLM can naturally generate patches that only pass the test containing "src" without considering patch robustness. This scenario is similar to the overfitting problem in APR [65]. To check for the overfitting problem, two of the authors independently evaluate the patch correctness, followed by a third person to review and confirm their disagreement to ensure the accuracy of the validation. In our manual validation, we also found that when D4C can generate five or more plausible patches for a buggy program in 10 samplings, most of these patches are correct. This observation aligns with previous work (i.e., responses generated multiple times tend to be more reliable [66]). DebugBench does not suffer from the overfitting problem because the given test examples are not used in its unseen test suite when checking for patch correctness.
As we allow D4C to debug (instead of relying on fault localization), we also observe that some correct patches have been applied at different locations compared to the developer patches. Fig. 8 shows an example where the developer patch modifies two if-conditions so it is a 2-hunk bug in the dataset. However, D4C only needs to insert one line, which combines the two conditions and results in a correct patch that is semantically equivalent to the developer patch. Furthermore, if we ask GPT-4 to generate patches at these two hunks, GPT-4 can only write two Exception statements, and fails to pass the unit tests. This example shows that allowing LLMs to generate patches without restricting the location may provide more flexibility, allowing generation of correct patches.
VI Conclusion
This paper presents a new approach to adapt LLMs to automated program repair. Our key insight is that the effectiveness of LLM-based APR can be enhanced by (1) aligning the output to their training objective and (2) allowing them to refine the whole program without first performing fault localization. Based on this insight, we designed D4C, an LLM-based prompting framework for APR. Our evaluation shows that D4C outperforms the SOTA APR methods with perfect fault localization by 10%. Our findings call for adopting a new paradigm and debugging workflow for future LLM-based APR approaches.